OpenStudy (karolynxx):

What is the value of x in the equation 3x –1/9 y = 18, when y = 27? a)5 b)7 c)45 d)63

OpenStudy (karolynxx):

@freckles help?

OpenStudy (freckles):

have you replace y with 27?

OpenStudy (freckles):

also that says: \[3x-\frac{1}{9}y=18\]

OpenStudy (freckles):

\[3x-\frac{1}{9}(27)=18 \text{ since } y=27\] we solve for x now

OpenStudy (freckles):

let me know if you do not know how to do that and I can give you the first step

OpenStudy (karolynxx):

so i do 1/9 x 27

OpenStudy (freckles):

also known as 27/9

OpenStudy (karolynxx):

27/9 = 3

OpenStudy (freckles):

yes

OpenStudy (freckles):

3x-3=18

OpenStudy (karolynxx):

So do i plug that in somewhere

OpenStudy (freckles):

you want to isolate the 3x term do this by adding 3 on both sides 3x-3=18 3x-3+3=18+3 3x+0=18+3 3x=21

OpenStudy (freckles):

now there is one last step

OpenStudy (karolynxx):

21/3?

OpenStudy (freckles):

yep

OpenStudy (karolynxx):

7!

OpenStudy (freckles):

yes

OpenStudy (freckles):

actually just 7 but I know what you mean 7! actually means 7*6*5*4*3*2*1 :p

OpenStudy (karolynxx):

Lol :-)

OpenStudy (karolynxx):

What is the value of x in the equation 7x + 2y = 48 , when y = 3?

OpenStudy (freckles):

so first step replace y with 3

OpenStudy (karolynxx):

7x + 2(3) =48 &x + 6=48

OpenStudy (freckles):

right 7x+6=48

OpenStudy (karolynxx):

so then id - 6 on both sides?

OpenStudy (freckles):

then two more steps

OpenStudy (freckles):

yes

OpenStudy (freckles):

one more after

OpenStudy (karolynxx):

so 42 then divode by 7?

OpenStudy (freckles):

yep

OpenStudy (freckles):

And karol I sleep now

OpenStudy (freckles):

good luck on your math

OpenStudy (karolynxx):

Thanks!

OpenStudy (freckles):

welcome :)